OpenStudy (anonymous):

24x²y / 40xy² reduces to

OpenStudy (shamil98):

\[\huge \frac{ 24 \times x \times x \times y }{ 40 \times x \times y \times y }\] What terms can you cancel, and what terms can you reduce?

OpenStudy (anonymous):

i got 3x/5y

OpenStudy (anonymous):

thanks

OpenStudy (shamil98):

yeah you got it right